OpenStudy (anonymous):

Evaluate the summation of 25 times 0.3 to the n plus 1 power, from n equals 2 to 10.. @ankit042

OpenStudy (anonymous):

\[\sum_{2}^{10}25\left( .3 \right) ^{n+1}=25 \left( .3 \right)^2\frac{ 1-\left( .3 \right)9}{ 1-.3 }=?\]

OpenStudy (anonymous):

correction \[=25\left( .3 \right)^3\frac{ 1-\left( .3 \right)^9 }{ 1-.3 }\]